### v1

# End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI

Abstract
Reconstructing cardiac motion from cine CMR sequences is critical for diagnosis, prediction, and intervention. Existing methods rely on complete CMR stacks to infer full heart motion, limiting their utility in intra-procedural scenarios where only sparse observations are available. We present TetHeart, the first end-to-end framework that unifies full 4D multi-structure heart mesh recovery from both offline full-stack acquisitions and intra-procedural sparse-slice observations. Our method leverages deep deformable tetrahedra, an explicit–implicit hybrid representation, to capture shape and motion in a coherent space shared across cardiac structures. It is initialized from high-quality pre-procedural or offline-acquired full stacks to build detailed, patient-specific heart meshes, which can then be updated using whatever slices are available, from full stacks down to a single slice. We further incorporate several key innovations: (i) an attentive mechanism for slice-adaptive 2D–3D feature assembly that dynamically integrates information from arbitrary numbers of slices at any position, combined with a distillation strategy from full-slice to sparse-slice settings to ensure accurate reconstruction under extreme sparsity; and (ii) a two-stage weakly supervised motion learning scheme requiring only keyframe (e.g., ED and ES) annotations. Trained and validated on three large public datasets and externally evaluated zero-shot on additional private interventional and public CMR datasets, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
{IEEEkeywords}
cardiac magnetic resonance (CMR), deformable tetrahedra, explicit–implicit hybrid representation, shape and motion modeling, sparse observation.
1 Introduction
The heart is one of the most vital organs in the human body [1], and accurate modeling of its shape and motion is central to numerous medical applications ranging from diagnosis to therapy planning [2, 3, 4]. Cardiac magnetic resonance (CMR) offers rich temporal imaging of cardiac anatomy and has become an indispensable modality for both shape [5, 6, 7, 8, 9] and motion [10, 11, 12, 13, 14] modeling. However, to this day, such modeling remains confined to offline analysis and pre-operative planning. Existing techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] cannot be applied during an intervention, as they require full-stack CMR data that cannot be acquired in real time. In the intervention room, only a handful of real-time slices can be captured, usually at lower resolution [15, 16, 17], a limitation also confirmed in routine practice at our collaborating university hospital. This scenario motivates us to move beyond the current state of the art and recover full 3D shape and motion from sparse and partial observations. This will enable real-time tracking of patient-specific heart dynamics, with a view to delivering guidance to interventional teams.
To this end, we introduce TetHeart, a unified framework for reconstructing the full 4D motion of multiple cardiac structures from either complete CMR stacks or intra-procedural sparse data. As in Fig. 1, TetHeart first leverages high-quality pre-operative or offline-acquired CMR stacks to build a detailed, patient-specific heart mesh, and then updates it in an end-to-end manner using any available slices, from a complete stack down to a single slice. To make this possible, TetHeart builds on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes, forming the foundation for jointly modeling shape and motion in a coherent space shared across structures. On top of this, our first core contribution is an attentive mechanism for slice-adaptive 2D-3D feature assembly, which enables the model to handle sparse input by dynamically integrating information from arbitrary slices at arbitrary positions. Additionally, we introduce a distillation strategy from full-slice to sparse-slice settings, ensuring accurate reconstruction under extreme input sparsity. Our second contribution is a two-stage weakly supervised motion learning scheme that requires only keyframe (e.g., ED and ES) annotations, making it feasible to train TetHeart across diverse clinical environments. The core idea is to first train a static shape reconstruction branch on labeled frames, and then repurpose it to initialize shapes for unlabeled frames, from which motion dynamics are recovered on the labeled frames to provide supervision.
Figure 1: TetHeart reconstructs cardiac motion from observed 2D slices. At $t=0$, a 3D CMR stack $\mathbf{I}^{0}$ is processed by the AFA module with self-attention to extract a volumetric static feature $V_{\text{static}}^{0}$, which is then used to generate the initial tetrahedra $\mathcal{G}^{0}$. For each subsequent time step $t$, the observed 2D slices $\mathbf{O}^{t}$ are encoded via the AFA module with cross-attention to obtain a volumetric motion feature $V_{\text{motion}}^{t}$, which is then used to recover motion by predicting offsets that deform $\mathcal{G}^{0}$. Weights are shared between the static reconstruction branch and dynamic deformation branch.
By representing the heart in a shared deformable space, TetHeart preserves spatial alignment between structures while adapting to varying sparsity levels. This offline/online capability enables comprehensive, patient-specific motion tracking across the entire clinical workflow: not only achieving state-of-the-art accuracy for offline use with full-stack CMR, but also, more importantly, unlocking online use in intra-procedural scenarios where only sparse observations are available.
We conducted extensive experiments to train and validate our model on three large public datasets: ACDC [19], M&Ms [20], and M&Ms-2 [21]. Remarkably, Remarkably, with a single set of weights learned from only end-diastolic (ED) and end-systolic (ES) annotations, our method generalizes seamlessly across both online few-slice and offline full-slice settings, consistently outperforming existing approaches not only on geometric metrics (Chamfer Distance and Dice) but also on clinically relevant indexes related to ejection fraction. We further performed external evaluations without retraining to assess generalizability. First, we tested the method on a privately collected dataset tailored for the online interventional MRI scenario at our collaborating cardiac CMR center, demonstrating its potential for deployment in real-world clinical environments. In addition, we evaluated the model on the public 4DM dataset [14], which provides complete temporal annotations of the myocardium, thereby demonstrating that the proposed weakly supervised motion learning scheme enables accurate reconstruction across the entire cardiac cycle. Together, these results highlight the effectiveness, robustness, and clinical promise of TetHeart. We will make our code public.
2 Related Work
2.1 Static Cardiac Shape Modeling
Reconstructing the heart from 3D medical images holds significant clinical value. However, cardiac MR (CMR) images often exhibit low through-plane resolution, which makes accurate reconstruction a challenge. Over the years, many approaches to addressing it have been proposed [22, 5, 23, 24, 25, 26, 27, 28, 29, 9] . They can be categorized into traditional optimization-based techniques and deep learning-based methods.
2.1.1 Optimization-Based Approaches
They usually rely on a two-stage pipeline. Contours are first extracted from the CMR images—either manually or using a segmentation algorithm—and then a template mesh is optimized to fit these contours. For example, the approach of [23] starts from a predefined tubular initial mesh that is deformed to minimize a contour-matching loss. More recently, a 3D active shape model (3D-ASM) and an image intensity model was used in [29] to align initial meshes with ground-truth contours. While these approaches can produce good results, they also have drawbacks. As they rely on test-time optimization, the reconstruction process is often time-consuming, taking from tens of seconds to several minutes per mesh. When applied to a full image sequence, this may extend to over an hour, making them impractical for real-time applications. Moreover, these methods’ performance has only been evaluated using ground-truth contours, which require labor-intensive manual annotation. What their performance is when using predicted contours instead remains an open question.
2.1.2 Deep Learning-Based Approaches
Deep learning-based methods leverage prior knowledge learned from large datasets to accelerate mesh reconstruction. For instance, in [28], sets of contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions. Similarly, in [26], contour points are embedded into a 3D volume and a hybrid architecture made of 3D CNNs and GCNs is used to progressively deform a template mesh. These methods increase speed but also suffer from several limitations. Because meshes are reconstructed solely from extracted contours or points, the rich appearance information contained in cine images is not fully exploited. Furthermore, as contour extraction is not necessarily perfect, such two-states approaches can result in error accumulation [26].
2.2 Dynamic Cardiac Motion Modeling
Arguably, when reconstructing deforming hearts from sequences of CMR scans, one could simply run a static reconstruction algorithm on each individual frame. However, this would fail to exploit temporal consistency and make it difficult to establish inter-frame correspondences. Thus a number of motion-based approaches that compute the deformation from frame to frame have been proposed.
2.2.1 Motion from Full-Slice Offline MRI Sequences
When full CMR stacks are available, recent deep learning-based approaches [11, 12, 13, 14, 6] can be used. MulViMotion [12] trains 3D CNNs to estimate left ventricular myocardial motion by predicting voxel-based deformation fields from CMR data, and then building 3D meshes by warping segmentation masks. DeepMesh [13] decouples mesh reconstruction from motion estimation, and uses 3D CNNs and bilinear interpolation to predict per-vertex deformation. 4DMR [14] models the myocardium using implicit representation. Decoupled shape and motion latent codes are optimized during test time to predict an initial shape and the backward motion fields. We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best currently available.
Unfortunately, these approaches also suffer from several limitations. First, some of them focus only on myocardium, neglecting other cardiac structures, preventing comprehensive heart assessment. Second, optimization-based methods like 4DMR are time-consuming at inference time. And, finally and most importantly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
2.2.2 Motion from Sparse Online MRI Sequences
CMR imaging has a reputation for being a ”slow” modality. Acquiring a full MRI stack involves sequential slice acquisition over a 4–6 heartbeat breath-hold, which can take several minutes [16, 17]. While this is feasible for pre-interventional planning, it is not suitable for intra-procedural settings.
Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without the need for gating, synchronization, or repetition [16]. Its applications include complex electrophysiology procedures [30], congenital heart disease [31], and other cardiac interventions [15, 32]. RT-MRI provides excellent soft-tissue contrast for assessing interventions like ablations [33, 17] and allows for continuous visualization and tracking of devices, and it is free of ionizing radiation [30, 34].
Nevertheless, although RT-MRI enables rapid and continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [16]. Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) to be used during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [15, 32, 34, 16, 17]. We are not aware of any current method that can reconstruct 3D cardiac motion using this kind of data. Yet, delivering a high-quality, dynamic 3D heart model to the intervention room based on real-time observations is crucial for guiding high-precision minimally invasive cardiac procedures [35, 36, 15] and for real-time localization of devices within the heart [30, 37]. This is the application that motivated us to begin this study.
3 Method
In this section, we introduce TetHeart, our framework for accurate 4D mesh reconstruction from either full CMR scans or sparse sets of 2D slices, as depicted by Fig 1. We first introduce our heart presentation that is based on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes. We then introduce our Attentive 2D–3D Feature Assembler (AFA) module, which enables effective feature encoding from 2D slices with varying numbers and locations. Finally, we present our weakly supervised training scheme that makes it possible to train given only a limited number of annotations.
3.1 Formalization
To model deforming hearts, we use the deep deformable tetrahedra [18] formalism. We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF value, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDF values in $\mathcal{T}$, the set of all tetrahedrons. The explicit surface can then be extracted using the differentiable marching tetrahedra algorithm [38].
Given an initial 3D heart model $\mathcal{G}^{0}$ reconstructed from the end-diastolic (ED) frame $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$, our goal is to estimate a deformed model $\mathcal{G}^{t}$ at each time instant $t$, based on subsequent observations $\mathbf{O}^{t}$ that reflect the ongoing cardiac motion. We do this by updating the vertices of $\mathcal{G}^{0}$. We write
$\mathcal{G}^{t}=(\{\mathbf{v}_{i}^{t}\},\{s_{i}\},\mathcal{T})\;,\mathbf{v}_{i}^{t}=\mathcal{D}(\mathbf{v}_{i}^{0},\mathbf{O}^{t})\;,$ (1)
where $\mathbf{v}_{i}^{0}$ is a grid vertex from $\mathcal{G}^{0}$, $\mathbf{v}_{i}^{t}$ its counterpart after deformation. $\mathcal{D}$ is the deformation model which we instantiate in the following sections and makes it possible to recover deformations across the whole sequence.
The observations $\mathbf{O}^{t}$ can be obtained in two manners.
  1. 1.
    Online. For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as necessary. However, it has to be possible to acquire $\mathbf{O}^{t}$ at sufficient frame rate and, consequently, it can only comprise a small set—from 1 to 3—of 2D slices at lower spatial resolution and at arbitrary locations [15].
  2. 2.
    Offline. Full-slice CMR images are available both initially and at time $t$, which is acquired from a slow slice-by-slice scanning process. In this case, $\mathbf{O}^{t}$ and $\mathbf{I}^{0}$ are of the same dimension $\mathbb{R}^{D\times H\times W}$.
In the following sections, we first introduce the AFA module and then describe how it is incorporated in our approach to generate the initial shape model from $\mathbf{I}^{0}$ and then deforming it given $\mathbf{O}^{t}$. Whether it is a full scan or a sparse set of slices, we use the same approach in both cases.
3.2 Attentive 2D-3D Feature Assembler
Because the observations can be either a full CMR stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases. To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that relies on an attention mechanism that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the total number of slices. With this notation, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$. A full-slice observation is simply one where $S=D$.
In practice, we first perform 2D convolutions on each individual slice to create a collection of 2D feature maps $F^{0}_{2d}$ and $F^{t}_{2d}$ from $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$. Given that $\mathbf{O}^{t}$ is not necessarily in the same coordinate system as $\mathbf{I}^{0}$, we use the metadata associated to these slices in the DICOM header to position them spatially in the $\mathbf{I}^{0}$ volume. This links each 2D location in the feature maps to a precise 3D location. In the remainder of this section, we first discuss how we use these 2D feature maps to compute volumetric motion features from $F^{0}_{2d}$ and $F^{t}_{2d}$, which are obtained from slices acquired at different times and can be used to estimate deformations. We then describe how we derive static features from a dense set of slices all acquired at the same time, which can be used to reconstruct the initial static model.
Motion Features. A naive way to use an attention mechanism to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values. However, this would be inefficient because global attention aims at capturing long-range dependencies, which is not needed to model the inherently local nature of anatomical motion. Also, this requires computing full pairwise relationships between all query and key positions, which is computationally expensive when using many slices and high-resolution feature maps.
To avoid this, we take our inspiration from the Swin Transformer [39] and use instead a variant of attention that preserves locality and significantly reducing computational complexity. For each voxel in $F^{0}_{2d}$, we first identify $k_{1}$ spatially nearest 2D slices, or all slices if there are only $k_{1}$ or less. Within each selected slice, we then retrieve the features at the $k_{2}$ spatially closest positions to serve as keys and values. Attention is subsequently computed only between the query and its corresponding set of localized keys and values. This can be written as
$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (2)
$\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{t}_{2d})\;,$
where NN-Select denotes the nearest slice and position selecting operation. MHAttention denotes a Multi-Head Attention [40] operation. A learnable position embedding is used to encode the spatial information into the calculation of attention. In practice, we set $k_{1}=3$ and $k_{2}=9$. This configuration ensures that the computational complexity of our attention mechanism is approximately equivalent to that of a 3D convolution with a kernel size of 3. As a result, the AFA module achieves high computational efficiency while retaining the flexibility and adaptability of attention-based feature aggregation.
While our AFA module effectively encodes features from varying 2D slice configurations, a persistent challenge remains in the online scenario, where the limited number of slices can hinder accurate 3D motion reconstruction. To alleviate this issue, we introduce a distillation loss that encourages the model to efficiently transfer knowledge from full-slice inputs to partial-slice inputs during training. In short, we predict two versions of motion features from full-slice observation and its randomly sampled subset. The distillation loss encourages the encoded feature from partial slices to be close to the feature from full-slice input. We formulate the loss we minimize to achieve this in Section 3.5.
Static Features. The motion features described above relate $\mathbf{I}^{0}$ to $\mathbf{O}^{t}$. The same mechanism can also be used to relate $\mathbf{I}^{0}$ to itself and produce features that can be used for static reconstruction. This can be similarly written as
$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (3)
$\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{0}_{2d})\;.$
Generating Final Features. After obtaining the volumetric static and motion features $V^{0}_{\rm static}$ and $V^{t}_{\rm motion}$, we use a U-Net-like [41, 42] architecture for further encoding. Just as standard U-Net architecture, it also consists of a downsampling and an upsampling stream. Features from the last level of the upsample stream are taken as the final features, which we denoted as $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$, respectively.
Note that the initial 2D encoding block and the subsequent 3D convolutions are shared for both static and motion feature extraction. This parameter sharing strategy enabled by our unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task. Consequently, it facilitates more accurate motion estimation—an advantage overlooked by some previous works [13, 14]. As demonstrated in the experiments, this architecture leads to superior motion reconstruction performance.
3.3 Initial Heart Model Reconstruction
Given the $F^{0}_{\rm static}$ features computed as discussed above from a full-slice CMR image $\mathbf{I}^{0}$, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$. To this end, we start from an initial tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube. We then trilinearly interpolate $F^{0}_{\rm static}$ at the vertices of $\mathcal{G}^{init}$. An SDF value and an offset for each vertex is then predicted using a GCN to create $\mathcal{G}^{0}$. We write
$\mathcal{G}^{0}$ $=(\{\mathbf{v}_{i}^{\prime}\},\{s_{i}\},T),$
$\mathbf{v}_{i}^{\prime}$ $=\mathbf{v}_{i}+\Delta\mathbf{v}_{i},$ (4)
$(\Delta\mathbf{v}_{i},s_{i})$ $=\operatorname{GCN}([F^{0}_{\rm static}(\mathbf{v}_{i}),\mathbf{v}_{i}])\;,$
where $\mathbf{v}_{i}$ is a grid vertex of $\mathcal{G}^{init}$ and $\mathbf{v}_{i}^{\prime}$ the same grid vertex after deformation. $F^{0}_{\rm static}(\mathbf{v})$ denotes feature extraction at $\mathbf{v}$ from $F^{0}_{\rm static}$ with trilinear interpolation, and $\cdot$$\cdot$] denotes concatenation. To enhance the network’s spatial sensitivity, the vertex location is appended to the feature vector before it is passed to the GCN. For datasets with $C$ classes, we simply set the output channel of the GCN layer to $C\times 4$, 3 for deformation, and 1 for SDF value. This yields a separate tetrahedral mesh for each class. Following prior works [43, 44, 45, 46], our network also has an auxiliary segmentation prediction branch during training time.
3.4 Motion Recovery
The static reconstruction scheme described above delivers comparable or even better reconstruction performance to that of approaches based on pure 3D U-Net architectures [43, 44, 45, 46], with the added benefit that it also allows the independent encoding of individual 2D slices. We now exploit this mechanism to compute deformations for the initial static shape from sets of slices acquired later. Given $\mathcal{G}^{0}$, $F^{0}_{\rm static}$, and $F^{t}_{\rm motion}$ computed as described above, we use a combination of GCN and GRU [47] layers to effectively aggregate spatial information and instantiate the deformation model $\cal{D}$ of Eq. 1. We write
$F^{cat}(\mathbf{v}_{i}^{(s)})$ $=[F^{0}_{\text{static}}(\mathbf{v}_{i}^{(s)}),F^{t}_{\text{motion}}(\mathbf{v}_{i}^{(s)})]\;,$ (5)
$F^{gcn}(\mathbf{v}_{i}^{(s)})$ $=\operatorname{GCN}(F^{cat}(\mathbf{v}_{i}^{(s)}))\;,$ (6)
$\mathbf{h}_{i}^{(s+1)}$ $=\operatorname{GRU}([F^{gcn}(\mathbf{v}_{i}^{(s)}),\mathbf{v}_{i}^{(s)}],\mathbf{h}_{i}^{(s)})\;,$ (7)
$\mathbf{v}_{i}^{(s+1)}$ $=\mathbf{v}_{i}^{(s)}+\text{MLP}(\mathbf{v}_{i}^{(s)},\mathbf{h}_{i}^{(s+1)})\;.$ (8)
In Eq. 5, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at the vertex locations. They are then concatenated and passed through Eq. 6-8 to gradually update the vertex position, which aggregates the spatial information. Eq. 5-8 is repeated $K$ times, which we set to $2$ by default.
3.5 Two-Stage Weakly Supervised Training Using Only Keyframe Annotations
Creating high-quality annotations for an entire sequence is costly. As a result, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases. This rules out full supervision. To address this, we propose a two-stage weakly-supervised training scheme which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
3.5.1 Training the Reconstruction Branch
We start by training the reconstruction branch of Section 3.3. At each training iteration, we randomly select a labeled image $\mathbf{I}$ with ground truth segmentation and mesh annotation $L^{gt},\mathcal{M}^{gt}$. Let $L^{p},\mathcal{G}^{p}$ be our model’s segmentation and tetrahedra prediction. We minimize the loss
$\mathcal{L}_{shape}$ $(\mathbf{I})=\lambda_{cd}\mathcal{L}_{cd}(\textbf{MT}(\mathcal{G}^{p}),\mathcal{M}^{gt})$ (9)
$+\lambda_{sdf}\mathcal{L}_{sdf}(\mathcal{G}^{p},\mathcal{M}^{gt})+\lambda_{ce}\mathcal{L}_{ce}(L^{p},L^{gt})\;,$
where $\mathcal{L}_{cd}$ is the chamfer distance, $\mathcal{L}_{sdf}$ is an L1-loss used to supervise the predicted SDF values at tetrahedral grid vertices with the ground truth SDF value queried from the ground truth mesh. $\mathcal{L}_{ce}$ is the cross-entropy loss defined on the predicted segmentation map. MT denotes the marching tetrahedra algorithm [38] to convert $\mathcal{G}^{p}$ into surface mesh. $\lambda_{cd},\lambda_{sdf},\lambda_{ce}$ are set to $1.0,0.1,0.1$ respectively.
3.5.2 Training the Deformation Estimator
We then train the motion branch. At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and then choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it. We first predict $\mathcal{G}^{u}$ from $\mathbf{I}^{u}$. Next, assuming that $\mathbf{I}^{l}$ comprises a total of $D$ slices, we randomly pick 1 to $D$ 2D slices at random location to form $\hat{\mathbf{I}}^{l}$ to simulate the online few-slice case. $V_{\text{motion}}^{l}$ and $\hat{V}_{\text{motion}}^{l}$ are generated from $\mathbf{I}^{l}$ and $\hat{\mathbf{I}}^{l}$ using the AFA module. We write
$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$ (10)
We use these 3D features to generate two different versions of deformed tetrahedra using Eq. 1: $\mathcal{G}^{u\rightarrow l}$ from full-slice encoded feature $V_{\text{motion}}^{l}$, and $\hat{\mathcal{G}}^{u\rightarrow l}$ from few-slice encoded feature $\hat{V}_{\text{motion}}^{l}$. The network is trained by minimizing
$\mathcal{L}_{motion}$ $(\mathbf{I}^{l},\mathbf{I}^{u})=\mathcal{L}_{distill}(V_{\text{motion}}^{l},\hat{V}_{\text{motion}}^{l})$ (11)
$+\mathcal{L}_{cd}(\text{MT}(\mathcal{G}^{u\rightarrow l}),\mathcal{M}^{l})+\mathcal{L}_{cd}(\text{MT}(\hat{\mathcal{G}}^{u\rightarrow l}),\mathcal{M}^{l}),$
[…]
where $\mathcal{M}^{l}$ is the ground truth mesh for $\mathbf{I}^{l}$ and sg denotes the stop-gradient operation. Minimizing the distillation loss $\mathcal{L}_{distill}$ ensures that the 3D information from full-slice $\mathbf{I}^{l}$ is successfully distilled into the few-slice encoded feature, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map. As a result, minimizing $\mathcal{L}_{motion}$ guarantees good reconstruction results for both full and few-slice inputs.
Once trained, our model can be used for new patients without retraining. Note that our training strategy actually enables the model to predict cardiac deformation starting from any phase, not just the end-diastolic (ED) phase. Additionally, by employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations. As a result, even though the online scenario is simulated during training, our model demonstrates strong generalization capability. It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions despite these challenges as will be shown in the experiment section.
4 Experiments
4.1 Experimental Settings
Table 1: Few-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes. 1- / 5-Slice: Number of slices used to recover the motion. Due to space limit, we omit the standard deviation.
           1-Slice
Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
4DMR [14] SDF 29.53 - - 75.63 - - 19.87 - - 76.12 - - 19.37 - - 74.11 - -
MR-Net [26] Mesh 28.61 40.70 46.15 77.43 82.12 73.33 14.59 16.46 25.11 77.59 80.35 73.94 15.61 16.68 25.23 76.29 82.66 76.23
Ours-Mesh [44] Mesh 17.42 25.11 42.33 80.14 84.90 74.32 11.26 14.17 23.19 79.88 82.11 75.96 11.98 15.11 22.63 77.66 84.97 77.99
Ours-SDF [46] SDF 18.65 26.45 43.78 79.55 83.43 74.00 12.16 14.88 23.91 79.09 81.62 74.23 12.15 15.72 23.34 77.01 84.37 77.15
Ours Tet 15.24 23.65 40.64 82.25 86.20 75.42 9.76 12.37 21.42 81.00 84.45 77.32 10.12 13.89 20.27 79.33 86.06 79.65
5-Slice
Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
4DMR [14] SDF 24.56 - - 77.03 - - 16.16 - - 77.33 - - 15.52 - - 76.67 - -
MR-Net [26] Mesh 18.82 24.22 36.54 79.54 84.25 74.67 10.60 12.16 20.28 78.70 83.31 76.91 10.76 12.05 19.25 77.84 86.03 78.79
Ours-Mesh [44] Mesh 11.66 19.39 33.91 82.92 85.88 75.73 9.14 11.07 18.62 80.99 85.11 78.67 8.68 10.96 16.82 79.82 87.99 80.13
Ours-SDF [46] SDF 12.21 20.03 34.52 82.35 85.32 75.11 9.63 11.69 19.26 80.45 84.56 78.05 9.22 11.55 17.42 79.21 87.41 79.53
Ours Tet 10.22 18.31 32.61 84.38 87.32 77.23 7.93 9.72 17.39 82.83 86.76 79.27 7.30 9.96 15.58 81.24 89.37 81.48
Table 2: Clinical indexes evaluation on M&Ms dataset. 1- / 5- / Full-Slice denotes the number of slices used to recover motion.
  Method LVESV MAE $\downarrow$ LVEF MAE $\downarrow$ RVESV MAE $\downarrow$ RVEF MAE $\downarrow$
(ml) (%) (ml) (%)
1-Slice
MR-Net [26] 9.76 7.87 16.40 11.75
Ours-Mesh [44] 8.92 6.37 16.41 9.83
Ours-SDF [46] 9.35 6.82 17.04 10.21
Ours 7.51 4.83 15.03 8.51
5-Slice
MR-Net [26] 7.34 6.29 13.45 8.87
Ours-Mesh [44] 7.64 5.11 12.81 7.33
Ours-SDF [46] 7.93 5.32 13.12 7.59
Ours 6.86 4.30 11.88 6.55
Full-Slice
FFD [48] 23.53 14.95 24.06 15.87
dDemons [49] 33.97 22.11 63.03 38.52
MR-Net [26] 6.75 4.16 12.05 7.01
Ours-Mesh [44] 5.07 2.23 9.85 4.96
Ours-SDF [46] 5.22 2.34 9.98 5.01
Ours 4.84 2.04 9.65 4.77
4.1.1 Data
We train and validate our method on a unified dataset constructed from three large and diverse publicly available datasets: ACDC, M&Ms, and M&Ms-2 [19, 20, 21]. ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm. For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames. M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively. While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies. As in earlier studies [13, 14, 9, 3], we register and fit a Statistic Shape Model (SSM) [24, 5] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes. The unified dataset comprises 835 CMR sequences, randomly split into 501 (60%), 167 (20%), and 167 (20%) sequences for training, validation, and testing, respectively.
We use two additional datasets for external evaluation to assess generalization in a clinical context where model retraining is not feasible. The first dataset was collected at the Center for Interventional MRI, University Hospital Lausanne (CHUV), designed with the target online scenario in mind and hereafter referred to as CHUV. For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle. In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack. Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations. To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute. Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest. For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks. To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we further use the second external dataset, 4DM [14], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame. This dataset enables a rigorous evaluation of the proposed weakly supervised motion learning scheme.
4.1.2 Evaluation Metrics
For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame. We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality. Dice score is also reported to quantify the intra-slice segmentation results. Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes. LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
4.1.3 Implementation Details
In all experiments, the slices are resampled by linear interpolation to a spacing of $1.25\times 1.25$ mm. From this, the AFA module constructs 3D volume with an effective spacing of $1.25\times 1.25\times 2$ mm. The AFA module uses a 8-head multi-head attention with learnable 3D position embedding. In the reconstruction branch, a modified nnU-Net [42] is used as the image encoder. It comprises five downsampling and upsampling blocks with channel sizes [32, 64, 128, 256, 320]. We use the same GCN structure as in [18] with $3$ layers and $128$ channels. The initial tetrahedra resolution is $128$. For the motion branch, we use the identical encoding network structure as the reconstruction branch. For the deformation model, the hidden dimension of the GCN and GRU are both set to $128$. We use SGD optimizer with an initial learning rate of 0.01, a weight decay of 3e53e-5, and momentum of $0.99$. We train all models for $300/150$ epochs for the reconstruction/motion learning stage. All experiments are conducted using one V100 GPU.
4.1.4 Baselines
We compare our method against several existing approaches for 4D shape reconstruction. For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for the online few-slice setting. Therefore, we adapted several baselines so that they can operate under sparse inputs. 4DMR [14] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points. We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame. MR-Net [26] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations. Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper. Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
Figure 2: 1-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note that 4DMR can only output motion for myocardium.
Figure 3: Motion sequence prediction results using different number of slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
In addition, we include two modified baselines derived from shape reconstruction methods: MeshDeformNet [44] and DeepCSR [46]. To adapt them for motion reconstruction, we integrated into both the same nnU-Net backbone, AFA module, deformation model, and training pipeline as our method. Since their main differences with our full framework lie in the shape representation, we denote them as Ours-Mesh and Ours-SDF.
For the traditional offline setting, in addition to the above methods, we further compare against B-spline Free Form Deformation (FFD) [48], diffeomorphic Demons (dDemons) [49], MR-Net w. nnU-Net, MulViMotion and DeepMesh. FFD and dDemons are classical registration algorithms that have been widely used in recent cardiac motion tracking studies [50, 51, 52]. MR-Net w. nnU-Net adopts the MR-Net architecture but replaces ground-truth contours with those predicted by nnU-Net, featuring a two-stage automated pipeline. MulViMotion and DeepMesh are mesh-based approaches that predict deformation fields via 3D CNNs, but only for the myocardium class. For fair comparison, we adopt their SAX-view-only versions as mentioned in the original papers.
Unless explicitly stated otherwise, all methods are jointly trained on the unified dataset.
4.2 Motion Estimation from a Few Slices
Table 3: Full-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
  Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
FFD [48] Mesh 21.84 38.85 51.04 69.46 79.56 72.04 42.01 63.23 88.77 62.84 76.65 66.23 40.79 66.42 95.66 63.13 79.80 76.02
dDemons [49] Mesh 15.72 22.97 35.85 77.06 86.87 73.60 32.16 44.39 61.15 65.83 79.87 65.09 34.08 48.77 66.22 67.25 83.44 72.39
4DMR [14] SDF 10.89 - - 80.03 - - 8.42 - - 82.33 - - 8.25 - - 82.67 - -
MR-Net [26] Mesh 14.02 19.21 24.07 80.02 86.35 78.93 8.86 10.03 16.86 79.40 86.90 79.73 8.39 7.24 12.99 80.72 87.86 80.60
MR-Net w. nnU-Net Mesh 15.11 20.45 26.13 79.66 85.38 76.45 9.88 10.25 18.60 78.56 85.00 77.21 9.83 9.24 13.35 79.56 85.89 78.31
MulViMotion [12] Mesh 17.21 - - 78.10 - - 13.75 - - 72.30 - - 12.33 - - 68.21 - -
DeepMesh [13] Mesh 16.42 - - 79.89 - - 9.24 - - 78.51 - - 9.20 - - 77.75 - -
Ours-Mesh [44] Mesh 6.80 11.84 23.41 86.62 88.11 84.74 4.26 5.08 13.92 84.09 87.12 83.73 4.15 5.12 11.04 85.01 88.67 86.65
Ours-SDF [46] SDF 7.21 12.65 22.78 86.95 88.43 84.00 4.44 5.26 14.21 83.87 86.91 83.77 4.30 5.44 11.32 84.55 88.37 86.10
Ours Tet 6.63 11.51 22.15 87.12 88.96 84.21 4.11 4.85 13.68 84.31 87.43 83.99 4.02 5.04 10.82 85.33 89.13 86.94
Table 4: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.
  Method Training Dataset Myocardium CD (mm2) $\downarrow$
ACDC M&Ms M&Ms-2 4DM 1-Slice 5-Slice Full-Slice
FFD - - - - - - 11.51
dDemons - - - - - - 12.46
4DMR - 24.00 19.03 13.97
MR-Net - 15.87 11.56 7.97
DeepMesh - - - 8.32
Ours-Mesh - 13.31 10.32 5.97
Ours-SDF - 13.78 10.78 6.24
Ours - - - 25.43 19.42 13.81
- - - 16.91 12.54 8.23
- - - 17.88 13.03 8.83
- 11.33 9.08 5.51
Ours (Oracle) - - - 6.24 4.87 2.70
In this section we present results under the online few-slice scenario that has been the main motivation for this work. In Fig. 2, we show two representative motion reconstruction results using only the central slice as input: one from a normal (NOR) heart and another from a subject with dilated cardiomyopathy (DCM). Because of its stronger contractile function, the normal heart exhibits larger deformations than the DCM heart, making motion prediction more challenging, which is reflected by larger magnitudes in the error maps. Nevertheless, our method delivers accurate predictions in both cases. Furthermore, the segmentation results on slices at the apex and base location demonstrate that our model can provide reliable deformation estimation even in regions far from the slices used for motion inference. The complete reconstructed sequences together with the corresponding volume–time curves are shown in Fig. 3.
We report quantitative results on the whole test set in Tabs. 1 and at the top of Tab. 2. The first contains chamfer distances and dice scores and the second cardiac function index predictions. For consistency, we generated all these results using the central 1 or 5 slices even though slices at any position could be used, as shown in Section 4.4. TetHeart consistently outperforms our four baselines without requiring ground-truth segmentations or contours as input. Instead, it infers cardiac motion directly from raw CMR images without any time-consuming test-time optimization. As a result, it delivers fast inference—12 FPS on a single V100—with a potential for further acceleration by adopting a lightweight encoder architecture and leveraging model quantization techniques to meet the real-time constraints. Even when using a single slice, TetHeart delivers meaningful indices. All of this points toward its suitability for eventual clinical deployment.
Figure 4: Chamfer distance as a function of the number of slices used to capture the motion.
Although we incorporated into Ours-Mesh and Ours-SDF the same mechanisms for handling sparse features as in our complete method, the latter still delivers better results. This confirms that the tetrahedral representation we use is more effective at preserving spatial information within the object, making it a better choice for inferring 3D motion when only limited observations are available. To quantify this, we plot in Fig. 4 the performance of all three methods as a function of the number of slices used. The fewer, the greater the gap between our complete approach and the other two. Nevertheless, both variants still surpass prior approaches by a clear margin, demonstrating that design components such as the AFA module and our training strategy form a general and effective solution for handling incomplete slice scenarios.
The comparatively poor performance of 4DMR can be attributed to its reliance on a simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered in this study. Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs. The weaker performance of MR-Net likely stems from the fact that it reconstructs meshes solely from contour points, thereby neglecting the rich information available in cine images.
4.3 Motion Estimation from Full Stacks
Figure 5: Full-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note 4DMR and DeepMesh can only output for myocardium.
Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab. 3 and the bottom of Tab. 2. In Fig. 3 and 5, we show qualitative results on the same hearts as in Fig. 2, but now using all slices as input instead of only one.
Remarkably, even though the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice scenario. We attribute this to several factors. Consistent with the previous analysis, the relatively poor performance of MR-Net may be attributed to its neglect of the rich information available in cine images. Using predicted contours resulted in slightly inferior performance, indicating that such a two-stage automated pipeline is prone to error accumulation, a finding consistent with the conclusions in the original paper [26]. MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes. Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage informative features from the reconstruction branch. Moreover, our AFA module dynamically aggregates spatially adjacent features, further enhancing prediction quality.
4.4 External Evaluation: To Generalize without Retraining
We now test the generalizability of our model to the CHUV and 4DM datasets without retraining.
In Fig. 6, we show sequence predictions on CHUV with the spatial position of the real-time captured 2D slice visualized in the left. As can be seen, TetHeart reconstructs a plausible heart motion across the cardiac cycle under both rest and exercise situation. Moreover, the predicted volume–time curves further confirm the physiological plausibility of our motion reconstruction. In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, which aligns well with expectation.
As 4DM provides full-sequence annotations, but only for the myocardium. we evaluate the full-sequence prediction performance for the myocardium subclass, rather than focusing solely on the ES frame and report the results in Tab. 4. Here, check marks denote the dataset used for training. Oracle refers to training and testing directly on 4DM, serving as an upper bound for performance.
As shown in the table, our model demonstrates the best generalization abilities, achieving top performance in both few-slice and full-slice settings. In contrast, 4DMR exhibits poor generalization, likely due to significant distribution shifts across datasets and the limited representational capacity of its simple MLP architecture. While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind our method.
Moreover, training with a larger and more diverse dataset improves generalization. And consistent with the findings from previous evaluations, representing the heart with tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
Figure 6: Motion sequence prediction on the CHUV dataset. (Left) Spatial configuration of the initial SAX stack and the online observations. (Middle) Cardiac motion sequences for each subject under rest and exercise conditions. (Right) Volume–time curves.
4.5 Ablation Studies
Finally, we conduct ablation studies to evaluate our design choices. First, we examine the role of the AFA module. Simply stacking 2D features achieves comparable performance to AFA in the full-slice setting, but it fails to generalize to scenarios with variable numbers and positions of 2D slices. This highlights the necessity of a more flexible mechanism like AFA. Next, we evaluate the importance of spatial information in the AFA module. If spatial position encodings are removed from the attention computation, the model’s performance degrades moderately, indicating that spatial information is crucial for accurate feature aggregation. However, if we use only position encodings as the query and discard image features entirely, the performance drops substantially. This shows that both image features and spatial coordinates are essential and must work together for effective motion prediction. When training the motion branch, if we do not share parameters with the reconstruction branch and instead initialize the deformation branch randomly, the model performance drops significantly. This demonstrates the effectiveness of our parameter-sharing strategy: the 3D knowledge learned during reconstruction is successfully transferred to motion modeling.
Lastly, we ablate the impact of the distillation loss. Without it, the model achieves slightly better results in the full-slice setting, but the features encoded from sparse 2D inputs are less discriminative, leading to poorer performance under few-slice conditions. When the distillation loss is combined with our aggressive sampling strategy, the model achieves consistently stronger results across all settings, demonstrating the importance of this training scheme for improving robustness and generalization.
Table 5: Ablation study results on M&Ms dataset.
  Method M&Ms Myocardium CD (mm2) $\downarrow$
1-Slice 5-Slice Full-Slice
Ours 9.76 7.93 4.11
- AFA module - - 4.22
- position embedding 11.23 9.04 4.67
- image feature in query $Q$ 16.67 12.37 6.44
- shared encoding network 20.66 11.63 6.43
- distillation loss 10.34 8.12 4.03
5 Conclusion
We have proposed TetHeart, a unified framework for reconstructing heart motion from either full stacks of 2D MRI slices or from sparse sets, down to a single slice. We demonstrate, for the first time, that full 3D heart shape and motion can be inferred in sparse real-time CMR slices, making our approach suitable for guidance of interventional procedures. This breakthrough shifts cardiac reconstruction from retrospective analysis to live clinical use, laying the foundation for real-time digital twins of the beating heart. This capability not only has the potential to enable precise, image-guided interventions today, but also points to a future where digital twins of organs in motion guide therapies across medicine—adaptive, predictive, and personalized at the point of care.
[…]